# A tuning-free and scalable method for joint graphical model estimation with sharper bounds

### v2

@@ Abstract
Joint estimation of multiple graphical models (i.e., multiple precision matrices) has emerged as an important topic in statistics. Unlike separate estimation, joint estimation can leverage shared structural patterns across multiple graphs to yield more accurate results. In this paper, we present an efficient and tuning-free method named MIGHT (Multi-task Iterative Graphical Hard Thresholding) to jointly estimate multiple graphical models. We reformulate the joint model into a series of multi-task learning problems in a column-by-column manner, and then solve these problems by using an iterative algorithm based on the hard thresholding operator. Theoretically, we derive the non-asymptotic error bound for our method. We prove that, under proper signal conditions, our method attains selection consistency and an improved error bound, and also exhibits asymptotic normality—properties rarely explored in existing joint graphical model estimation literature. The performance of our method is validated through numerical simulations and real data analysis of a cancer gene-expression RNA-seq dataset.
Key words : Asymptotic normality, Joint estimation, Multi-task learning, Nodewise regression, Precision matrix estimation, Selection consistency
@@ 1 Introduction
We consider the joint estimation of multiple Gaussian graphical models in this paper. For $p$ fixed variables, we collect observations drawn from Gaussian distributions across $K$ different classes or datasets, denoted as $X^{(1)}\in\mathbb{R}^{n_{1}\times p},\cdots,X^{(K)}\in\mathbb{R}^{n_{K}\times p}$. Although their corresponding precision matrices, i.e., the inverse covariance matrices, $\Theta^{(1)},\cdots,\Theta^{(K)}\in\mathbb{R}^{p\times p}$ may exhibit dataset-specific heterogeneity, they often share many common support structures (Guo et al., 2011). This structural commonality motivates a joint estimation of these graphical models instead of separate estimations.
Joint graphical models have significant practical implications in data analysis. For instance, in neuroimaging studies, by collecting $n$ fMRI scans from each of the $K$ participants, we might estimate the functional connectivity network (i.e., graphical model) among the $p$ voxels for each participant $k$. While the networks among participants have varying structures exhibiting heterogeneity, the shared common structures indicate functional correspondence. Leveraging these shared patterns, we can get improved estimation through joint modeling approaches (Tsai et al., 2022; van den Bergh et al., 2024). Additionally, transfer learning of graphical models has become a heated topic recently. In this context, joint estimation, which identifies shared structural patterns across graphical models, can serve as a pre-processing step for transfer learning to assess the similarity between source studies and the target study (Li et al., 2023; Zhao et al., 2024). Moreover, joint graphical model estimation plays a significant role in gene expression analysis (Lee and Liu, 2015), stock price analysis (Yang and Peng, 2020), graph neural networks (Jiang et al., 2022), and so on.
In this paper, we propose an efficient and adaptive method for the joint graphical model estimation. Through theoretical analysis and practical applications, we demonstrate the advantages of our method in parameter estimation and inference.
@@ 1.1 Related works
Joint graphical model estimation has garnered significant attention over the past decade due to its ability to leverage shared structural patterns across multiple datasets. Pioneering work like Guo et al. (2011) introduced hierarchical regularization to decompose and identify the common shared and class-specific edges among these precision matrices. Danaher et al. (2014) proposed two regularization methods: the Group Graphical Lasso (GGL) and the Fused Graphical Lasso (FGL). GGL combines an element-wise sparsity penalty, $\sum_{i\neq j}\sum_{k\in[K]}|\Theta_{ij}^{(k)}|$, with a covariate-wise penalty, $\sum_{i\neq j}\left\{\sum_{k\in[K]}\left(\Theta_{ij}^{(k)}\right)^{2}\right\}^{1/2}$, to encourage shared sparsity patterns among different graphical models. In contrast, FGL combines the same element-wise sparsity penalty with a fused-type penalty, $\sum_{k<k^{\prime}}\sum_{i\neq j}\left|\Theta_{ij}^{(k)}-\Theta_{ij}^{(k^{\prime})}\right|$, to encourage shared edge values among different graphical models. These two foundational methods have inspired lots of advanced joint estimation methods. Price et al. (2015) proposed a ridge penalty plus a ridge fused penalty to maintain the similarity among the multi-graphs. Gibberd and Nelson (2017) proposed the group-fused graphical lasso method to estimate the time-evolving and piecewise-constant Gaussian graphical model. Bilgrau et al. (2020) extended the work of Price et al. (2015) to the case with prior information. And Navarro et al. (2022) proposed a convex optimization method with $\ell_{1}$ fused penalty under the general case of graph stationarity. Additionally, joint estimation methods based on the CLIME estimator (proposed by Cai et al. (2011)) (Lee and Liu, 2015; Qiu et al., 2015; Cai et al., 2016), and Bayesian frameworks (Li et al., 2019; Jalali et al., 2023), have also been investigated. For a comprehensive review of joint graphical model estimation, we refer the reader to Tsai et al. (2022).
@@
While the aforementioned joint estimation methods capture shared structures, some critical challenges remain: (i) high computational cost, (ii) complex tuning procedures, and (iii) the lack of theoretical guarantees for selection consistency and inference. To address these issues, we begin by considering a group-structured linear model
@@
$y=\sum_{j=1}^{p}X_{G_{j}}\bm{\beta}_{G_{j}}+\epsilon,\quad\text{with }\underbrace{\sum_{j=1}^{p}\mathbf{1}(\bm{\beta}_{G_{j}}\neq\mathbf{0})\leq s}_{\text{group-wise sparsity}}\text{ and }\underbrace{\sum_{j=1}^{p}\|\bm{\beta}_{G_{j}}\|_{0}\leq ss_{0}}_{\text{element-wise sparsity}},$ (1)
@@
with the covariates partitioned into $p$ groups $\{G_{j}\}_{j=1}^{p}$. Within this joint sparse framework, Simon et al. (2013) introduced the sparse group lasso, which simultaneously applies both an element-wise penalty $\|\bm{\beta}\|_{1}=\sum_{j=1}^{p}\|\bm{\beta}_{G_{j}}\|_{1}$ and a group-wise penalty $\sum_{j=1}^{p}\|\bm{\beta}_{G_{j}}\|_{2}$. The penalty terms of the sparse group lasso and the GGL exhibit similar patterns, reflecting the analogous multiple sparsity structures in their underlying statistical models. Recently, there have been many cutting-edge works on the joint sparse model (1): Cai et al. (2022) provided theoretical guarantees for parameter estimation of the sparse group lasso method. Zhang et al. (2024) developed a double sparse iterative hard thresholding procedure to achieve the minimax adaptivity. Abramovich (2024) proposed a classifier based on logistic regression that leverages such joint structure (1). Motivated by these developments, we introduce a novel analytical framework that reformulates the joint graphical model into a series of linear models via the node-wise regression approach (Meinshausen and Bühlmann, 2006; Janková and van de Geer, 2017). Then, by developing enhanced techniques, this strategy can obtain more refined estimation and have the potential to overcome the limitations of current joint graphical estimation methods.
@@ 1.2 Our contributions
The main contributions of this work can be summarized in the following three aspects:
  • We propose a computationally efficient method called MIGHT (Multi-task Iterative Graphical Hard Thresholding) to jointly estimate multiple graphical models with common structures. We reformulate the joint graphical model estimation problem as a series of multi-task learning problems, which are then solved in a column-by-column fashion. This transformation enables our method to support parallel computation, significantly enhancing its efficiency and scalability in practical applications.
  • We establish rigorous theoretical guarantees for the MIGHT method. We derive sharper estimation error bounds compared to existing methods and prove that MIGHT achieves selection consistency under proper signal conditions. Additionally, we demonstrate the asymptotic properties of MIGHT, a key feature for statistical inference that is typically absent in regularization-based methods.
  • We implement the MIGHT method in an open-source R package ADSIHT, available at https://cran.r-project.org/. The comprehensive numerical experiments on synthetic and real datasets demonstrate that our method achieves superior empirical performance than the state-of-the-art methods.
@@ 1.3 Organization and Notations
The current paper is organized as follows. Section 2 introduces our method and algorithm. Section 3 derives the theoretical guarantees of our method. Sections 4 and 5 demonstrate the advantages of our method through simulations and empirical data, respectively. Section 6 concludes this paper. All technical proofs and additional simulations are provided in the supplementary material.
@@ 2 Methods
This section introduces the model setup and transforms the joint graphical model into a series of multi-task learning problems. An iterative hard thresholding-type algorithm is then introduced to solve these problems.
@@ 2.1 Model setup
We consider the joint estimation of the graphical models across $K$ independent datasets. For the $k$-th dataset, observations are organized into an $n_{k}\times p$ matrix $X^{(k)}=\left(x_{1}^{(k)},\cdots,x_{n_{k}}^{(k)}\right)^{\top}$, where each row $x_{\ell}^{(k)}$ follows an independent $p$-dimensional Gaussian distribution $N(\bm{\mu}^{(k)},\Sigma^{(k)})$, for every $\ell\in[n_{k}]$. Without loss of generality, we assume $\bm{\mu}^{(k)}=\mathbf{0}_{p}$. Denote by $\Theta^{(k)}:=\left(\Sigma^{(k)}\right)^{-1}\in\mathbb{R}^{p\times p}$ the underlying precision matrix of the $k$-th dataset. For $i\neq j$, the entry $\Theta_{i,j}^{(k)}$ quantifies the conditional dependence between nodes $i$ and $j$ in the $k$-th graph, with $\Theta_{i,j}^{(k)}=0$ indicating conditional independence of $i$ and $j$ given the remaining $p-2$ covariates (Lauritzen, 1996). We denote by $S_{j}^{(k)}$ the neighbor set of node $j$ in the $k$-th graphical model, i.e.,
@@
and let $S_{j}:=\bigcup_{k=1}^{K}S_{j}^{(k)}$ denote the total neighbor set of node $j$ across all $K$ graphs, which includes nodes that connect to node $j$ in at least one graphical model. Define $N:=\sum_{k\in[K]}n_{k}$.
Figure 1: The joint graphical models with $K=4$ graphs and $p=8$ nodes. (a) The edges in these graphs can be divided into three categories: edges universally shared across all graphs (the black lines), edges partially shared among subsets (the green lines), and edges unique to individual graphs (the red lines). (b) The corresponding adjacency matrices for each graph in (a). We take node 1 as an example, its neighbor sets across the four graphs are $S_{1}^{(1)}=\{4,5,7\},S_{1}^{(2)}=\{2,4,5\},S_{1}^{(3)}=\{4,5\}$, and $S_{1}^{(4)}=\{5,7\}$. The total neighborhood set $S_{1}=\bigcup_{k=1}^{4}S_{1}^{(k)}=\{2,4,5,7\}$ includes all nodes that connect to node 1 in at least one graph. On average, each edge that is connected to node $1$ appears in $\frac{\sum_{k=1}^{4}|S_{1}^{(k)}|}{|S_{1}|}=2.5$ graphs, reflecting the degree of similarity among these four graphs.
@@
We assume that these $K$ graphical models share some common structures, allowing joint estimation to achieve higher accuracy and lower computational cost than separate estimation (Guo et al., 2011; Danaher et al., 2014; Tsai et al., 2022). As illustrated in Figure 1, the joint graphical model exhibits two structural properties: neighbor-similarity and edge-sparsity, which are formalized in the following.
Definition 1 (Neighbor-similarity)
We say the joint graphical model $\left(\Theta^{(1)},\cdots,\Theta^{(K)}\right)$ is $s_{0}$ neighbor-similar ($1\leq s_{0}\leq K$) if
$\max_{j\in[p]}\frac{\sum_{k=1}^{K}|S_{j}^{(k)}|}{|S_{j}|}\leq s_{0}.$
We quantify the neighbor-similarity by constraining the average frequency of each neighbor, for every node $j\in[p]$. The parameter $s_{0}$ measures the similarity among the $K$ graphs: an $s_{0}$ approaching $K$ allows the near-complete neighbor overlap across all graphs, indicating homogeneous structural supports; in contrast, a smaller $s_{0}$ encourages each graph to possess some individual structure, demonstrating greater heterogeneity among these graphs.
Definition 2 (Edge-sparsity)
@@
We say the joint graphical model $\left(\Theta^{(1)},\cdots,\Theta^{(K)}\right)$ is $s$ edge-sparse ($10\leq s\leq p-$ 1) if
$\max_{j\in[p]}|S_{j}|\leq s.$
@@
We quantify the edge-sparsity by constraining the size of the total neighbor set for each node. It also provides the sparsity in each individual graphical model (since $|S_{j}^{(k)}|\leq|S_{j}|$ for each $(j,k)\in[p]\times[K]$). Below, we assume that the joint graphical model exhibits both $s_{0}$ neighbor-similarity and $s$ edge-sparsity, which we denote as the joint $(s_{0},s)$ structure.
@@
We next estimate the precision matrices in a column-by-column fashion, i.e., estimating all the $j$-th columns $\left(\Theta_{\cdot,j}^{(1)},\cdots,\Theta_{\cdot,j}^{(K)}\right)$ at once. This is inspired by the node-wise regression (Meinshausen and Bühlmann, 2006), which reformulates the estimation of a column in a precision matrix as a sparse linear regression problem (Janková and van de Geer, 2017) For each $(j,k)\in[p]\times[K]$, let $X_{j}^{(k)}\in\mathbb{R}^{n_{k}\times 1}$ denote the observation vector corresponding to the $j$-th covariate, and let $X_{\setminus j}^{(k)}\in\mathbb{R}^{n_{k}\times(p-1)}$ denote the submatrix of $X^{(k)}\in\mathbb{R}^{n_{k}\times p}$ excluding the $j$-th covariate. The next proposition indicates that the column-wise estimation is equivalent to a multi-task learning model.
@@
$\underbrace{\begin{pmatrix}X_{j}^{(1)}\\ X_{j}^{(2)}\\ \vdots\\ X_{j}^{(K)}\end{pmatrix}}_{=:X_{j}\in\mathbb{R}^{N}}=\underbrace{\begin{pmatrix}X_{\setminus j}^{(1)}&\mathbf{O}_{n_{1}\times(p-1)}&\cdots&\mathbf{O}_{n_{1}\times(p-1)}\\ \mathbf{O}_{n_{2}\times(p-1)}&X_{\setminus j}^{(2)}&\cdots&\mathbf{O}_{n_{2}\times(p-1)}\\ \vdots&\vdots&\ddots&\vdots\\ \mathbf{O}_{n_{K}\times(p-1)}&\mathbf{O}_{n_{K}\times(p-1)}&\cdots&X_{\setminus j}^{(K)}\end{pmatrix}}_{=:X_{\setminus j}\in\mathbb{R}^{N\times(p-1)K}}\underbrace{\begin{pmatrix}\bm{\alpha}_{\setminus j}^{(1)}\\ \bm{\alpha}_{\setminus j}^{(2)}\\ \vdots\\ \bm{\alpha}_{\setminus j}^{(K)}\end{pmatrix}}_{=:\bm{\alpha}_{\setminus j}\in\mathbb{R}^{(p-1)K}}+\underbrace{\begin{pmatrix}\epsilon_{j}^{(1)}\\ \epsilon_{j}^{(2)}\\ \vdots\\ \epsilon_{j}^{(K)}\end{pmatrix}}_{=:\epsilon_{j}\in\mathbb{R}^{N}},$ (2)
where $N=\sum_{k\in[K]}n_{k}$, and $\epsilon_{j}^{(k)}\in\mathbb{R}^{n_{k}}$ is independent of $X_{\setminus j}^{(k)}$, and
$\bm{\alpha}_{\setminus j}^{(k)}=-\left(\Theta_{j,j}^{(k)}\right)^{-1}\Theta_{\setminus j,j}^{(k)}\in\mathbb{R}^{p-1},\quad\epsilon_{j}^{(k)}\sim N\left(\mathbf{0}_{n_{k}},\left(\Theta_{j,j}^{(k)}\right)^{-1}\cdot\mathbf{I}_{n_{k}}\right).$ (3)
Since $S_{j}^{(k)}=\text{supp}\left(\Theta_{\setminus j,j}^{(k)}\right)=\text{supp}\left(\bm{\alpha}_{\setminus j}^{(k)}\right)$, after the column-wise transformation (2), the joint $(s_{0},s)$ structure of the joint graphical model is encoded to the coefficient vector $\bm{\alpha}_{\setminus j}$, for each $j\in[p]$. Therefore, our main objective is to estimate the joint precision matrices $\left(\Theta^{(1)},\cdots,\Theta^{(K)}\right)$ column-by-column. By leveraging both neighbor-similarity and edge-sparsity structures, we aim for an accurate joint estimation measured by the weighted squared column-wise $\ell_{2}$ norm $\sum_{k=1}^{K}\frac{n_{k}}{N}\left\|\hat{\Theta}_{\cdot,j}^{(k)}-\Theta_{\cdot,j}^{(k)}\right\|_{2}^{2}$.
Additionally, to facilitate subsequent algorithmic analysis, we apply column-wise scaling to $X_{\setminus j}$ in model (2), ensuring that each column has an $L_{2}$-norm of order $\sqrt{N}$. The scaled model is then expressed as
@@
where $\bm{\beta}_{\setminus j}\in\mathbb{R}^{(p-1)K}$ is composed of $\bm{\beta}_{\setminus j}^{(1)},\cdots,\bm{\beta}_{\setminus j}^{(K)}\in\mathbb{R}^{p-1}$, see Algorithm 2 for specific rescale operations. The scaling operations preserve the support structure, i.e., $\text{supp}\left(\bm{\beta}_{\setminus j}^{(k)}\right)=\text{supp}\left(\bm{\alpha}_{\setminus j}^{(k)}\right)=S_{j}^{(k)}$.

### v3

# Learning Joint Graphical Models with Computational Efficiency, Dynamic Regularization, and Adaptation

@@ Abstract
Multi-sourced datasets are common in studies of variable interactions, for example, individual-level fMRI integration, cross-domain recommendation, etc, where each source induces a related but distinct dependency structure. Joint learning of multiple graphical models (i.e., multiple precision matrices) has emerged as an important tool in analyzing such data. Unlike separate learning, joint learning can leverage shared structural patterns across graphs to yield more accurate results. In this paper, we present an efficient and adaptive method named MIGHT (Multi-task Iterative Graphical Hard Thresholding) to estimate multiple graphs jointly. We reformulate the joint model into a series of multi-task learning problems through a column-by-column manner, and solve these problems using a dynamic regularized algorithm based on iterative hard thresholding. This framework is inherently parallelizable and therefore efficient in computation. Theoretically, we derive the non-asymptotic error bound for the resulting estimator. Furthermore, the proposed algorithm is adaptive to heterogeneous column-wise signal strengths: for nodes with strong signals, our estimator achieves improved error bounds and selection consistency adaptively, and also exhibits asymptotic normality—properties rarely explored in existing joint learning methods. The performance of our method is illustrated through numerical simulations and real data analysis on a cancer gene-expression RNA-seq dataset.
Key words : Graphical model, Joint estimation, Multi-task learning, Nodewise regression, Precision matrix estimation, Selection consistency
@@ 1 Introduction
This paper analyzes the joint learning of multiple Gaussian graphical models. For $p$ fixed variables, assume that we collect observations drawn from Gaussian distributions across $K$ different classes or datasets, denoted as $X^{(1)}\in\mathbb{R}^{n_{1}\times p},\cdots,X^{(K)}\in\mathbb{R}^{n_{K}\times p}$. Although their corresponding precision matrices, i.e., the inverse covariance matrices, $\Theta^{(1)},\cdots,\Theta^{(K)}\in\mathbb{R}^{p\times p}$ may exhibit dataset-specific heterogeneity, they often share many common support structures Guo et al. (2011). This structural commonality motivates a joint estimation of these graphical models, rather than separate estimations.
Joint graphical models have significant practical implications in data analysis. For instance, in neuroimaging studies, by collecting $n$ fMRI scans from each of the $K$ participants, we can estimate the functional connectivity network (i.e., graphical model) among the $p$ voxels for each participant. The networks among participants have some shared common structures, indicating functional correspondence. Leveraging these patterns, we can get improved estimation through joint modeling approaches Tsai et al. (2022); van den Bergh et al. (2025). Additionally, transfer learning of graphical models has become a heated topic recently. In this context, joint estimation, which identifies shared structural patterns, can serve as a pre-processing step for transfer learning to assess the similarity between source studies and the target study Li et al. (2023); Zhao et al. (2024). Moreover, joint graphical model estimation plays a significant role in gene expression analysis Lee and Liu (2015), stock price analysis Yang and Peng (2020), graph neural networks Jiang et al. (2022), and so on. These findings necessitate the study of joint graph estimation.
In this paper, we propose an efficient and adaptive method for joint graphical model learning. Through theoretical analysis and practical applications, we demonstrate the advantages of our method in statistical inference.
@@ 1.1 Related works
Joint graphical model  Joint graphical model estimation has garnered significant attention over the past decade due to its ability to leverage shared structural patterns across multiple datasets. Pioneering work like Guo et al. (2011) introduced hierarchical regularization to decompose and identify the common shared and class-specific edges among these precision matrices. Danaher et al. (2014) proposed two regularization methods: the Group Graphical Lasso (GGL) and the Fused Graphical Lasso (FGL). GGL combines an element-wise sparsity penalty, $\sum_{i\neq j}\sum_{k\in[K]}|\Theta_{ij}^{(k)}|$, with a covariate-wise penalty, $\sum_{i\neq j}\left\{\sum_{k\in[K]}\left(\Theta_{ij}^{(k)}\right)^{2}\right\}^{1/2}$, to encourage shared sparsity patterns among different graphical models. In contrast, FGL combines the same element-wise sparsity penalty with a fused-type penalty, $\sum_{k<k^{\prime}}\sum_{i\neq j}\left|\Theta_{ij}^{(k)}-\Theta_{ij}^{(k^{\prime})}\right|$, to encourage shared edge values among different graphical models. These two foundational methods have inspired lots of advanced joint estimation methods. Price et al. (2015) proposed a ridge penalty plus a ridge fused penalty to maintain the similarity among the multi-graphs. Gibberd and Nelson (2017) proposed the group-fused graphical lasso method to estimate the time-evolving and piecewise-constant Gaussian graphical model. Bilgrau et al. (2020) extended the work of Price et al. (2015) to the case with prior information. And Navarro et al. (2022) proposed a convex optimization method with $\ell_{1}$ fused penalty under the general case of graph stationarity. Additionally, joint estimation methods based on the CLIME estimator (proposed by Cai et al. (2011)) Lee and Liu (2015); Qiu et al. (2015); Cai et al. (2016), and Bayesian frameworks Li et al. (2019); Jalali et al. (2023), have also been investigated. For a comprehensive review, we refer the reader to Tsai et al. (2022).
@@
Motivation and inspiration   While the aforementioned joint estimation methods capture shared structures, some critical challenges remain: (i) high computational cost, (ii) complex tuning procedures, and (iii) the lack of theoretical guarantees for selection consistency and inference. To address these issues, we begin by considering a group-structured linear model
@@
$y$ $=\sum_{j=1}^{p}X_{G_{j}}\bm{\beta}_{G_{j}}+\epsilon,\qquad\text{with }\underbrace{\sum_{j=1}^{p}\mathbf{1}(\bm{\beta}_{G_{j}}\neq\mathbf{0})\leq s}_{\text{group-wise sparsity}}\text{ and }\underbrace{\sum_{j=1}^{p}\|\bm{\beta}_{G_{j}}\|_{0}\leq ss_{0}}_{\text{element-wise sparsity}},$ (1)
@@
with the covariates partitioned into $p$ groups $\{G_{j}\}_{j=1}^{p}$. Within this joint sparse framework, Simon et al. (2013) introduced the sparse group lasso, which simultaneously applies both an element-wise penalty $\|\bm{\beta}\|_{1}=\sum_{j=1}^{p}\|\bm{\beta}_{G_{j}}\|_{1}$ and a group-wise penalty $\sum_{j=1}^{p}\|\bm{\beta}_{G_{j}}\|_{2}$. The penalty terms of the sparse group lasso and the GGL exhibit similar patterns, reflecting the analogous multiple sparsity structures in their underlying statistical models. Recently, there have been many cutting-edge works on the joint sparse model (1): Cai et al. (2022) provided theoretical guarantees for parameter estimation of the sparse group lasso method. Zhang et al. (2024) developed a double sparse iterative hard thresholding procedure to achieve the minimax adaptivity. Abramovich (2024) proposed a classifier based on logistic regression that leverages such a joint structure (1). Tugnait (2021, 2024) extended the sparse-group-lasso-type penalty to learning multi-attribute graphical models and matrix-valued graphical models, respectively. Motivated by these developments, we introduce an analytical framework that reformulates the joint graphical model into a series of linear models via the nodewise regression approach Meinshausen and Bühlmann (2006); Janková and van de Geer (2017). Then, by developing a dynamic regularized procedure, we can obtain a more refined estimation while maintaining computational efficiency. Our method overcomes the limitations of current joint graphical estimation methods.
@@ 1.2 Our contributions
The main contributions of this work are threefold:
  • We propose a computationally efficient method called MIGHT (Multi-task Iterative Graphical Hard Thresholding) to jointly estimate multiple graphs with common structures. We reformulate the joint graphical model estimation problem as a series of multi-task learning problems, and then solve them using a dynamic regularized algorithm based on the iterative hard thresholding. This transformation enables parallel computation, enhancing its efficiency and scalability in practical applications.
  • We establish theoretical guarantees for the MIGHT method. We derive sharper estimation error bounds compared to existing methods and prove that MIGHT achieves selection consistency under proper signal conditions. Additionally, we demonstrate the asymptotic properties of MIGHT, a key feature for statistical inference that is typically absent in regularization-based method literature.
  • We implement the MIGHT method in an open-source R package ADSIHT, available at https://cran.r-project.org/web/packages/ADSIHT/index.html. The comprehensive numerical experiments on synthetic and real datasets demonstrate that our method achieves superior empirical performance than the state-of-the-art methods.
@@ 1.3 Organization and Notations
The current paper is organized as follows. Section 2 introduces our method and algorithm. Section 3 derives the theoretical guarantees of our method. Sections 4 and 5 demonstrate the advantages of our method through simulations and empirical data, respectively. Section 6 concludes this paper. All technical proofs and additional simulation results are provided in the appendices and the supplementary material.
@@ 2 Methods
This section introduces the model setup and transforms the joint graphical model into a series of multi-task learning problems. A dynamic-regularized iterative hard thresholding-type algorithm is then introduced to solve these problems.
@@ 2.1 Model setup
Consider the joint estimation of the graphical models across $K$ independent datasets. For the $k$-th dataset, observations are organized into an $n_{k}\times p$ matrix $X^{(k)}=\left(x_{1}^{(k)},\cdots,x_{n_{k}}^{(k)}\right)^{\top}$, where each row $x_{\ell}^{(k)}$ follows an independent $p$-dimensional Gaussian distribution $N(\bm{\mu}^{(k)},\Sigma^{(k)})$, for every $\ell\in[n_{k}]$. Without loss of generality, we assume $\bm{\mu}^{(k)}=\mathbf{0}_{p}$. Denote by $\Theta^{(k)}:=\left(\Sigma^{(k)}\right)^{-1}\in\mathbb{R}^{p\times p}$ the underlying precision matrix of the $k$-th dataset. For $i\neq j$, the entry $\Theta_{i,j}^{(k)}$ quantifies the conditional dependence between nodes $i$ and $j$ in the $k$-th graph, with $\Theta_{i,j}^{(k)}=0$ indicating conditional independence of $i$ and $j$ given the remaining p2p-2 covariates Lauritzen (1996). We denote by $S_{j}^{(k)}$ the neighbor set of node $j$ in the $k$-th graphical model, i.e.,
@@
and let $S_{j}:=\bigcup_{k=1}^{K}S_{j}^{(k)}$ denote the total neighbor set of node $j$ across all $K$ graphs, which includes nodes that connect to node $j$ in at least one graph. Define $N:=\sum_{k\in[K]}n_{k}$. We assume that these $K$ graphical models share some common structures, allowing joint estimation to achieve higher accuracy and lower computational cost than separate estimation Guo et al. (2011); Danaher et al. (2014); Tsai et al. (2022). As illustrated in Figure 1, the joint graphs exhibit two structural properties: local edge-sparsity and local neighbor-similarity, which are formalized in the following.
Figure 1: The joint graphical models with $K=4$ graphs and $p=8$ nodes. (a) Edges in these graphs can be divided into three categories: edges universally shared across all graphs (black lines), edges partially shared among subsets (green lines), and edges unique to individual graphs (red lines). (b) Corresponding adjacency matrices for (a). Take node 1 as an example, its neighbor sets across the four graphs are $S_{1}^{(1)}=\{4,5,7\},S_{1}^{(2)}=\{2,4,5\},S_{1}^{(3)}=\{4,5\}$, and $S_{1}^{(4)}=\{5,7\}$. The total neighborhood set $S_{1}=\bigcup_{k=1}^{4}S_{1}^{(k)}=\{2,4,5,7\}$ includes all nodes that connect to node 1 in at least one graph. On average, each edge that is connected to node 11 appears in $\frac{\sum_{k=1}^{4}|S_{1}^{(k)}|}{|S_{1}|}=2.5$ graphs, reflecting the degree of (local) similarity among these graphs.
Definition 1 (Local edge-sparsity)
@@
We say the joint graphical model $\left(\Theta^{(1)},\cdots,\Theta^{(K)}\right)$ is local edge-sparse with parameters $\{s_{j}\}_{j\in[p]}$ if
$|S_{j}|=s_{j},\quad\text{for every }j\in[p].$
@@
We quantify the local edge-sparsity by constraining the total neighbor set for each node $j$ respectively (e.g., $s_{1}=4$ in Figure 1). It also provides a sparse neighbor set in each individual graphical model (since $|S_{j}^{(k)}|\leq|S_{j}|$ for each $(j,k)\in[p]\times[K]$).
Definition 2 (Local neighbor-similarity)
We say the joint graphical model $\left(\Theta^{(1)},\cdots,\Theta^{(K)}\right)$ is local neighbor-similar with parameters $\{s_{0,j}\}_{j\in[p]}$ if
$\frac{\sum_{k=1}^{K}|S_{j}^{(k)}|}{|S_{j}|}=s_{0,j},\quad\text{for every }j\in[p].$
We quantify the local neighbor-similarity by constraining the average frequency of each neighbor for every node $j\in[p]$. An $s_{0,j}$ approaching $K$ allows the near-complete neighbor overlap across all graphs, indicating locally homogeneous structure for node $j$. In contrast, a smaller $s_{0,j}$ encourages each graph to possess some individual neighbors to node $j$, demonstrating greater local heterogeneity among these graphs.
@@
Assume that the joint graphical model exhibits both local edge-sparsity (with $\{s_{j}\}_{j\in[p]}$) and local neighbor-similarity (with $\{s_{0,j}\}_{j\in[p]}$), which we denote as the joint $(s_{j},s_{0,j})_{j\in[p]}$ structure. We next estimate the precision matrices in a column-by-column fashion, i.e., estimating all the $j$-th columns $\left(\Theta_{\cdot,j}^{(1)},\cdots,\Theta_{\cdot,j}^{(K)}\right)$ at once. This is inspired by the nodewise regression Meinshausen and Bühlmann (2006), which reformulates the estimation of a column in a precision matrix as a sparse linear regression problem Janková and van de Geer (2017). For each $(j,k)\in[p]\times[K]$, let $X_{j}^{(k)}\in\mathbb{R}^{n_{k}\times 1}$ denote the observation vector corresponding to the $j$-th covariate, and let $X_{\setminus j}^{(k)}\in\mathbb{R}^{n_{k}\times(p-1)}$ denote the submatrix of $X^{(k)}\in\mathbb{R}^{n_{k}\times p}$ excluding the $j$-th covariate. The next proposition indicates that the column-wise estimation is equivalent to a multi-task learning model.
@@
$\underbrace{\begin{pmatrix}X_{j}^{(1)}\\ \vdots\\ X_{j}^{(K)}\end{pmatrix}}_{=:X_{j}\in\mathbb{R}^{N}}=$ $\underbrace{\begin{pmatrix}X_{\setminus j}^{(1)}&\cdots&\mathbf{O}_{n_{1}\times(p-1)}\\ \vdots&\ddots&\vdots\\ \mathbf{O}_{n_{K}\times(p-1)}&\cdots&X_{\setminus j}^{(K)}\end{pmatrix}}_{=:X_{\setminus j}\in\mathbb{R}^{N\times(p-1)K}}\underbrace{\begin{pmatrix}\bm{\alpha}_{\setminus j}^{(1)}\\ \vdots\\ \bm{\alpha}_{\setminus j}^{(K)}\end{pmatrix}}_{=:\bm{\alpha}_{\setminus j}\in\mathbb{R}^{(p-1)K}}+\underbrace{\begin{pmatrix}\epsilon_{j}^{(1)}\\ \vdots\\ \epsilon_{j}^{(K)}\end{pmatrix}}_{=:\epsilon_{j}\in\mathbb{R}^{N}},$ (2)
where $N=\sum_{k\in[K]}n_{k}$, and
$\bm{\alpha}_{\setminus j}^{(k)}$ $=-\left(\Theta_{j,j}^{(k)}\right)^{-1}\Theta_{\setminus j,j}^{(k)}\in\mathbb{R}^{p-1},$ (3)
$\epsilon_{j}^{(k)}$ $=X_{j}^{(k)}-X_{\setminus j}^{(k)}\bm{\alpha}_{\setminus j}^{(k)}\sim N\left(\mathbf{0}_{n_{k}},\left(\Theta_{j,j}^{(k)}\right)^{-1}\cdot\mathbf{I}_{n_{k}}\right).$
Since $S_{j}^{(k)}=\text{supp}\left(\Theta_{\setminus j,j}^{(k)}\right)=\text{supp}\left(\bm{\alpha}_{\setminus j}^{(k)}\right)$, after the column-wise transformation (2), the joint $(s_{j},s_{0,j})_{j\in[p]}$ structure is encoded to the coefficient vectors $\left\{\bm{\alpha}_{\setminus j}\right\}_{j\in[p]}$. Additionally, to facilitate subsequent algorithmic analysis, we apply column-wise scaling in model (2), ensuring that each column of $X_{\setminus j}$ has an $L_{2}$-norm of order $\sqrt{N}$. The scaled model is then expressed as
@@
where $\bm{\beta}_{\setminus j}\in\mathbb{R}^{(p-1)K}$ is composed of $\bm{\beta}_{\setminus j}^{(1)},\cdots,\bm{\beta}_{\setminus j}^{(K)}\in\mathbb{R}^{p-1}$, see Algorithm 2 for specific rescale operations. Therefore, our main objective is to estimate the precision matrices $\left(\Theta^{(1)},\cdots,\Theta^{(K)}\right)$ in a column-by-column fashion, leveraging both local neighbor-similarity and edge-sparsity to achieve accurate and computationally efficient joint estimation.